If f(*)=$(1/*^2 +2* -3)d* find f(*) given that f(2) =2


1
Expert's answer
2021-01-19T12:36:55-0500

"If\ f(x)=(\frac{1}{x^2} +2x -3)dx.\\ Find\ f(x)\ given\ that\ f(2) =2"


This means "f'(x)=x^{-2}+2x-3"


Integrate "f'(x)" :

"f'(x)=x^{-2}+2x-3,\intop\frac{1}{x^2}dx+\intop 2xdx-\intop 3dx=-\frac{1}{x}+x^2-3x+c\\
f(x)=\frac{1}{x}+x^2-3x+c,\\
But\ f(2)=2\\
f(2)=2=\frac{1}{2}+2^2-3(2)+c,\\
c=\frac{9}{2}\\
\implies f(x)=\frac{1}{x}+x^2-3x+\frac{9}{2}\\"